16) how to solve 4x=y-1 with the substitution of 6x-2y=-3
Answer by mananth(12270)

how to solve 4x=y-1 with the substitution of 6x-2y=-3
..
6x-2y=-3
6x=2y-3
/6
x=y/3-1/2
...
4(y/3-1/2)= y-1
(4y/3)-2=y-1
(4y/3)-y = 2-1
(4y-3y)/3=1
y/3= 1
y=3
plug balue of y in
4x=y-1
4x=3-1
4x=2
x=1/2
...
m.ananth@hotmail.ca